OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0. Enter exponents using the caret (^). 9-5x^2=(x-2)^2-4

OpenStudy (anonymous):

Why is this in history?

OpenStudy (anonymous):

some dude needed help in history and i helped him

OpenStudy (anonymous):

Lolz anyways.... \[9-5x^2=(x-2)^2-4\]\[9-5 x^2 = x^2-4 x\]Then move it over \[-6 x^2+4 x+9 = 0\]

OpenStudy (anonymous):

Also, move to Mathematics